OpenStudy (anonymous):

What is the vertex and the equation of the axis of symmetry of the graph of y=x^2-6x-7?

OpenStudy (zzr0ck3r):

for parabola \(f(x) = ax^2+bx+c\) the vertex is \[(\frac{-b}{2a},f(\frac{-b}{2a}))\] and the axis of symmetry is \(x=\frac{-b}{2a}\)

OpenStudy (anonymous):

Okay, I'm a bit confused. :o lol. Sorry, I'm terrible at math.

OpenStudy (anonymous):

Is it C?

OpenStudy (zzr0ck3r):

you have \(x^2-6x-7\) compare that to \(ax^2+bx+c\) what is a?

OpenStudy (anonymous):

Oh, sorry, forgot I didn't give you the options. a) vertex: (–16, 3) a.o.s: x = –16 b) vertex: (–3, 20) a.o.s: x = –3 c) vertex: (3,–16) a.o.s: x = 3 d) vertex: (20, –3) a.o.s: x = 20

OpenStudy (pinksapphire):

The correct answer is letter C.

OpenStudy (anonymous):

Thank you.

OpenStudy (pinksapphire):

Use a graphing Calculator to help you...

OpenStudy (zzr0ck3r):

just take the class for her....

OpenStudy (anonymous):

Lol. I already figured it was C, I just needed some confirmation. I understand the work though.